# EVF-SAM: Early Vision-Language Fusion for Text-Prompted Segment Anything Model

### v4

@@ 4.2 Implementation Details
Table 2: Comparison of cIoU on different benchmarks between our proposed EVF-SAM and previous state-of-the-art methods. Bold: the best results. Underline: the second-best results. AVG represents the average metric across the eight RefCOCO-series benchmarks. We abbreviate the datasets: COCO (C) (Lin et al., 2014), RefCOCO (RC) (Yu et al., 2016; Nagaraja et al., 2016; Mao et al., 2016; Kazemzadeh et al., 2014), Objects365 (O) (Shao et al., 2019), Video segmentation datasets (V), ADE20K (A) (Zhou et al., 2017a; 2019), COCO-Stuff (CS) (Caesar et al., 2018), PACO-LVIS (PL) (Ramanathan et al., 2023), PASCAL-Part (PP) (Chen et al., 2014), GranD (G) (Rasheed et al., 2023), PASCAL VOC2010 (PV) (Everingham et al., 2010), MUSE (M) (Ren et al., 2023), gRefCOCO (gRC) (Liu et al., 2023a), COCO-Interactive (CI) (Zhang et al., 2024), FSS-1000 (F) (Li et al., 2020), SA-1B (SA) (Kirillov et al., 2023), PartImageNet (PIN) (He et al., 2022), HumanParsing (HP) (Liang et al., 2015b; a), GoldG (GG) (Kamath et al., 2021).
@@
EVF-SAM BEIT-3 RC 82.1 83.7 80.0 75.2 78.3 70.1 76.8 77.4 78.0
EVF-SAM BEIT-3 RC, O, A, PP, PIN, HP 82.4 84.2 80.2 76.5 80.0 71.9 78.2 78.3 79.0
4.3 Main Results
We mainly report the cIoU metric of RefCOCO-series benchmarks and compare our proposed EVF-SAM with recent state-of-the-art methods in Tab. 2 The upper part of Tab. 2 presents traditional methods based on text encoders. Despite their advantages in terms of fewer parameters and faster inference speeds, these methods either achieve less competitive results or require vast amounts of data due to their lack of integration with foundation models. The methods listed in the lower portion of Tab. 2 are based on Large Multimodal Models (LMMs), achieving state-of-the-art (SOTA) performance but require significant computational resources. Our EVF-SAM achieves the highest average cIoU score across all RES benchmarks, using only limited data and manageable computation costs. Specifically, our EVF-SAM achieves SOTA performance on RefCOCOg (Nagaraja et al., 2016; Mao et al., 2016), predicating a stronger capability for handling longer text prompts than previous LMM-based models, which is counter-intuitive while showing the great potential of vision-language models for understanding instructions. In addition, the early fusion between the input image and text prompts can generate more informative embeddings than independent encoders as discussed in Sec. 1.
4.4 Ablation Study
In this section, we conduct experiments to investigate the vision-language models for text-prompted SAM and study the effects of the designs of the proposed EVF-SAM. Unless specified, we mainly report the cIoU on testA of RefCOCO.
@@ 4.4 Ablation Study
Table 3: Ablation on fusion methods. We evaluate the performance of using different pre-trained Multimodal Encoders in EVF-SAM, e.g., CLIP from OpenAI (Radford et al., 2021) or OpenCLIP (Ilharco et al., 2021). $L_{i}$ denotes the $i$-th layer in the BEIT-3 model (totally 24 layers for BEIT-3-Large). Half of the layers are activated to assess the impact of the modality fusion stage on model performance. $\dagger$: pre-trained models provided by OpenAI. $\ddagger$: pre-trained models provided by OpenCLIP.
@@
Multimodal Encoder and fusion methods. In Tab. 3, we explore the effects of different Multimodal Encoders, e.g., CLIP, ViLT (Kim et al., 2021), and BEIT-3, and fusion methods, e.g., late fusion or early fusion. As shown in Tab. 3, using a text-only encoder in EVF-SAM obtains limited segmentation performance on RefCOCO. Using Multimodal Encoders with both image and text inputs remarkably improves 4.5 cIoU, 4.6 cIoU, 4.0 cIoU, 1.0 cIoU, and 4.5 cIoU for CLIP-Large (OpenAI  [^3] OpenAI: https://github.com/openai/CLIP), CLIP-Large (OpenCLIP  [^4] OpenCLIP: https://github.com/mlfoundations/open_clip), CLIP-Huge (OpenCLIP), ViLT, and BEIT-3, respectively. It demonstrates the superiority of using multimodal prompts (text and input image) and showcases that the image embeddings will also provide useful guidance for SAM to segment objects accurately. We further evaluate the effects of early fusion on ViLT and BEIT-3, which adopts modality fusions in all self-attention layers. Specifically, we adopt two settings for BEIT-3 to analyze, e.g., fusions among former 12 layers ($L_{1}\sim L_{12}$), and fusions among all layers ($L_{1}\sim L_{24}$). Tab. 3 indicates that BEIT-3 with early fusion (fusing former 12 layers or fusing all 24 layers) significantly improves compared to late fusion or using text only. In addition, ViLT with early fusion also achieves 11.1 cIoU improvements compared to the baseline with text-only prompts, showing the effectiveness of early fusion and multimodal inputs for prompting SAM. Therefore, Tab. 3 demonstrates that (1) Multimodal Encoder with the input image and text and (2) early fusions between the image and text encoder are much effective for text-prompted SAM.
Ablations on trainable modules. In Tab. 5, we evaluated the effects of fine-tuning (✓) or freezing ($\ast$) modules in the proposed EVF-SAM, i.e., the Multimodal Encoder, the prompt encoder, and the mask decoder. The image encoder of SAM is kept frozen during training. As Tab. 5 shows, fine-tuning the Multimodal Encoder is crucial and it adapts the Multimodal Encoder to encode text and image inputs to multimodal representation for referring image segmentation. Notably, EVF-SAM can achieve competitive results with all modules of SAM kept frozen, and it can be seamlessly regarded as a strong extension for the original SAM, which simultaneously supports text prompts, box prompts and point prompts. Tab. 5 Further fine-tuning the prompt encoder and mask decoder of SAM brings significant improvements.
Table 4: Ablations on trainable modules. We mainly evaluate the effects of fine-tuning or freezing the Multimodal Encoder, the prompt encoder and mask decoder of SAM. ‘✓’ denotes trainable, while $\ast$’ denotes frozen.
Multimodal Enc. Prompt Enc. Mask Dec. cIoU
$\ast$ $\ast$ 21.2
$\ast$ $\ast$ 82.9
$\ast$ 83.3
83.7
[…]
- 83.5
- 83.7
- 83.5
Concat 83.2
Multimodal feature representation. In Tab. 5, we explore the effects of using different multimodal features representations as prompts for SAM. Specifically, we adopt different outputs of the Multimodal Encoder: (a) the image [CLS] token, (b) the AvgPool over image tokens, and (c) the text [CLS] token. Tab. 5 shows that using image [CLS] token is more effective while combining image and text tokens through concatenation leads to a performance drop.
Effects of Different Foundation Models. In Tab. 6, we explore the effects of using different foundation models in EVF-SAM. For the Multimodal Encoder, we adopt CLIP-Large (only text encoder), ViLT, BEIT-3-Large, and BEIT-3-Base. We also modify EVF-SAM with Efficient-SAM (Xiong et al., 2023) to formulate a lighter version, which reduces 600M parameters compared to SAM-H. As shown in Tab. 6, EVF-SAM with BEIT-3-Base brings a severe performance drop which indicates a better Multimodal Encoder leads to better prompts for SAM. Remarkably, Tab. 6 shows a negligible difference between Efficient-SAM-S and SAM-H in EVF-SAM, which demonstrates the effectiveness of Efficient-SAM and also indicates that EVF-SAM performs well for different SAM variants. In addition, it also provides insights about designing text-prompted SAMs for future research, e.g., developing a larger and better Multimodal Encoder is more important to empower SAM with text-prompted abilities.
Table 6: Comparison of effects of different foundation models. AVG represents the average metric across the 8 RefCOCO-series benchmarks.
@@
4.5 Discussions

### v5

@@ 2.3 Late-fusion and Early-fusion
Table 1: Comparison of cIoU on different benchmarks between our proposed EVF-SAM and previous state-of-the-art methods. Bold: the best results. Underline: the second-best results. AVG represents the average metric across the eight RefCOCO-series benchmarks. We abbreviate the datasets: COCO (C) [33], RefCOCO (RC) [72, 42, 41, 21], Objects365 (O) [51], Video segmentation datasets (V), ADE20K (A) [77, 78], COCO-Stuff (CS) [3], PACO-LVIS (PL) [46], PASCAL-Part (PP) [5], GranD (G) [47], PASCAL VOC2010 (PV) [11], MUSE (M) [50], gRefCOCO (gRC) [35], COCO-Interactive (CI) [75], FSS-1000 (F) [28], SA-1B (SA) [24], PartImageNet (PIN) [12], HumanParsing (HP) [32, 31], GoldG (GG) [20].
@@
EVF-SAM BEIT-3 (673M) RC 82.1 83.7 80.0 75.2 78.3 70.1 76.8 77.4 78.0
EVF-SAM BEIT-3 (673M) RC, O, A, PP, PIN, HP 82.4 84.2 80.2 76.5 80.0 71.9 78.2 78.3 79.0
3 Method
3.1 Motivation: Early Vision-Language Fusion.
We investigate how to encode text prompts for SAM in this section. We start by using a text encoder following SAM’s trial, and extend to late-fusion and early-fusion, as shown in Fig. 3. We conduct preliminary experiments on RefCOCO (testA), shown in Tab. 2.
Table 2: Motivation analysis. Both CLIP and BEIT-3 are of Large scale, with comparable numbers of parameters. Specifically, CLIP has a total parameter count of 428M, while BEIT-3 totals 673M parameters. ‘L’ and ‘V’ denote the linguistic input and the vision input. The reported metrics are evaluated on RefCOCO/testA, except that the metric of LLaVA [36] is borrowed from LISA-7B [25].
CLIP (L) CLIP (L+V) BEIT-3 (L) BEIT-3 (L+V) LLaVA (L+V)
cIoU 63.4 67.9 65.1 83.7 79.1
Multimodal referring information is crucial. The interactive capability of SAM [24] hinges on its prompt encoder, which converts points, boxes, and masks into embeddings that incorporate positional information. As a result, when we are dealing with the Referring Expression Segmentation (RES) task based on SAM, the key challenge is to get the positional information from the input referring prompt. SAM [24] explored text-prompted segmentation with a CLIP text encoder, as illustrated in Fig. 3 (a). However, it is infeasible for the unimodal text encoder to extract positional information. Our reproduced metrics in Tab. 2 shows that CLIP-prompted SAM achieves 63.4 cIoU, which is inferior to well-defined baselines. Moreover, we observe performance improvements after using multimodal prompts, i.e., 63.4 v.s. 67.9 for CLIP and 65.1 v.s. 83.7 for BEIT-3, proving that multimodal referring information is effective for text-prompted SAM.
Early-fusion outperforms late-fusion. Existing works employ pre-encoded text feature (e.g., RES via text encoder) or pre-encoded image feature (e.g., RES via LLM), as illustrated in Sec. 2.2, demonstrating the ‘late-fusion’ manner. We visualize the attention map within multimodal fusion layers in Fig. 2, and find shortcomings within ‘late-fusion’ methods. The attention map W is calculated by $\textbf{W}=\text{clamp}_{0}^{1}(2\times\text{sigmoid}(QK^{\mathrm{T}}/\sqrt{d}))$, where $Q$ denotes query, $K$ denotes Key, and $\text{clamp}(\cdot)$ is a clipping function for a more distinct comparison We firstly observe enhanced image-to-text attention at the first few layers, indicating that ‘RES via text encoder’ methods, which lack image-to-text attention, may struggle with accurately interpreting text descriptions. Moreover, the ‘text-to-image attention’ takes the lead at later layers, showing the inefficiency of ‘RES via LLM’ methods, which results in a limited multimodal perception of visual features. The attention visualization example in Fig. 2 further illustrates the imprecise object localization capabilities of late-fusion methods. In contrast, our proposed ‘early-fusion’ methods which utilize shallow embedding layers and deep fusion layers (e.g., [23, 59]) is effective and efficient at both ‘image-to-text attention’ and ‘text-to-image attention’, overcoming the shortcomings of ‘late-fusion’. Experiment metrics in Tab. 3.1 further support our motivation. SAM with CLIP-text and CLIP-image, which implements a straightforward ‘late-fusion’ approach through concatenation, achieves a cIoU of 67.9. Furthermore, SAM with LLaVA, representing a more complex form of ‘late-fusion’, reaches 79.1 cIoU but requires 7B parameters. Conversely, SAM with BEIT-3, exemplifying our ”early-fusion” approach, achieves a cIoU of 83.7, surpassing all late-fusion methods while utilizing just 1.3 billion parameters.
3.2 Architecture
Multimodal Encoder. The Early Vision-Language Fused encoder adopts the input image and text and outputs fused multimodal embeddings. In EVF-SAM, we mainly adopt BEIT-3 [59] as the Multimodal Encoder, which formulates a multi-way transformer. The text is tokenized by XLMRobertaTokenizer [6] while the image is resized to $224^{2}$ and patched by a 1/16 convolution layer. Within each block of the encoder, the image and text tokens will be fused in the attention block and then fed into separate Feed-Forward Networks (FFN). We follow ViT [10] to retrieve the [CLS] token as the output multimodal embeddings.
Adapted prompt encoder for SAM. In EVF-SAM, we maintain the architecture of the image encoder and mask decoder of SAM, while extending the prompt encoder to further gather the embeddings from the Multimodal Encoder. Specifically, the original prompt encoder encodes point or box prompts to sparse embeddings of $\mathbb{R}^{B\times N\times D}$, where $B$ is the batch size, $N$ is the number of points/boxes, and $D$ is the embedding dimension. In EVF-SAM, the projected multimodal embeddings of $\mathbb{R}^{B\times 1\times D}$ from the Multimodal Encoder will be concatenated to sparse embeddings (of $\mathbb{R}^{B\times 0\times D}$, as there is no points/boxes inputs) and then fed into the mask decoder.
3.3 Data strategy
Since publicly available open-source RES data is limited, a simple thought is that we use datasets of other segmentation sub-tasks. To be specific, we include Instance Segmentation (Objects365 [51]), Semantic Segmentation (ADE20K [77, 78]) and Part Segmentation (Pascal-Part [5], HumanParsing [32], PartImageNet [12]) to boost the performance of our EVF-SAM, as well as to extend the model’s capability granularity.
However, segmentation data from different tasks exhibit varying distributions, making it challenging to effectively integrate them for joint training. We carefully analyze the semantic conflict between semantic segmentation data and RES data, and the ambiguity problem between instance segmentation data and RES data in Sec. A.1. Furthermore, we detailedly explain our data strategies in Sec. A.2, including applying a special token ‘[semantic]’ for solving semantic conflict, and using filtering/merging principles to solve the ambiguity problem.
Powered by our data strategies, we successfully integrate segmentation data from diverse tasks into an unified hybrid dataset. Upon joint training with the hybrid dataset, we observe the performance gain of EVF-SAM from average 78.0 cIoU to 79.0 cIoU, as shown in Tab. 1. Furthermore, EVF-SAM can handle various granularities of text-prompted segmentation such as semantic-level (e.g., stuff, multi-object) and part-level (e.g., body-part) segmentation. Detailed definition, metrics and visualizations are shown in Sec. A.3 and Sec. A.4, separately.
Table 3: Comparison of the model efficiency of EVF-SAM with LLM-based RES methods. Training time means the total time cost of training on RefCOCO [21, 72]. Evaluation time means the average time cost of forward during evaluation on refcoco/unc/testA.
Methods Training time Evaluation time cIoU
LISA-7B-LORA 54.7h 0.772s 76.5
EVF-SAM 19.7h
(-64%)
0.283s
(-63%)
84.2
(+10%)
Effi-EVF-SAM-S 8.2h
(-85%)
0.129s
(-83%)
83.5
(+9%)
4 Experiments
4.1 Datasets and Metrics
Datasets. We mainly conduct the experiments on RefCLEF [21], RefCOCO, RefCOCO+ [72, 21], and RefCOCOg [42, 41]. Specifically, RefCOCO+ excludes geometric expression (e.g., ‘on the left’), while RefCOCOg contains longer expressions (8 words per prompt on average). Among different splits of testing datasets, ‘testA’ is human-centric, while ‘testB’ aims for common objects.
To produce our best result, we employ multi-task unified training, as illustrated in Sec. 3.3. Objects365 [51], ADE20K [77, 78], Pascal-Part [5], HumanParsing [32] and PartImageNet [12] are additionally included.
4.2 Implementation Details
Unless specified, we initialize the proposed EVF-SAM with the public weights of SAM-ViT-Huge  [^1] SAM: https://github.com/facebookresearch/segment-anything [24] and BEIT-3-Large  [^2] BEIT-3: https://github.com/microsoft/unilm [59]. All models are trained on 4 NVIDIA L40s GPUs with mixed precision. We adopt DeepSpeed [53] with ZeRO-2 for model parallel to optimize memory consumption. During training, the batch size of each GPU is 16 and we use gradient accumulation for 2 steps, therefore, the total batch size per iteration is 128. We adopt AdamW [40] optimizer and set the initial learning rate to 1e-4 with a linear-decay schedule. We train all models for 15k iterations and use the binary cross-entropy loss (BCE) and dice loss (the weight of both losses is 1.0).
@@ 4.2 Implementation Details
Table 4: Ablation on fusion methods. We evaluate the performance of using different pre-trained Multimodal Encoders in EVF-SAM, e.g., CLIP from OpenAI [45] or OpenCLIP [17]. $L_{i}$ denotes the $i$-th layer in the BEIT-3 model (totally 24 layers for BEIT-3-Large). Half of the layers are activated to assess the impact of the modality fusion stage on model performance. $\dagger$: pre-trained models provided by OpenAI. $\ddagger$: pre-trained models provided by OpenCLIP.
@@
4.3 Main Results
We mainly report the cIoU metric of RefCOCO/+/g benchmarks and compare our proposed EVF-SAM with recent state-of-the-art methods in Tab. 1. The upper part of Tab. 1 presents traditional methods based on text encoders. Despite their advantages in terms of fewer parameters or task-specific architectures, these methods either achieve less competitive results or require vast amounts of data and computational resources due to their lack of integration with foundation models. The methods listed in the lower portion of Tab. 1 are based on Large Language Models (LLMs), achieving more competitive performance but at the expense of significantly increased parameter size. Our EVF-SAM achieves the highest average cIoU score across all RES benchmarks, using only readily available data, manageable computation resources, and moderate parameter size. Notably, unlike previous methods that excelled on specific datasets but performed poorly on others, our approach achieves consistently high scores across multiple datasets. This demonstrates that our method does not rely on specific linguistic expressions (e.g., geometric words) for prediction but rather possesses strong generalization capabilities.
Besides metrics on RefCOCO/+/g, we also evaluate the model’s performance on semantic-level referring expression segmentation, part-level referring expression segmentation and reasoning segmentation. Metrics are shown in Sec. A.3, Sec. A.4 and Sec. B.1.
4.4 Efficiency analysis
To prove the efficiency of our EVF-SAM over previous LLM-based models, we present comparisons of training and evaluation time in Tab. 3. We take LISA [25], which employs LLaVA [36] as the multimodal extractor, for comparison. We follow LISA’s reported setting to conduct 10k steps of training with a global batch of 160. Considering that different datasets contain descriptions of various lengths (e.g., 3.6 words per sentence for RefCOCO, 8.4 words per sentence for RefCOCOg), we conduct experiments only on RefCOCO to avoid sample bias. Our EVF-SAM can reduce 64% to 85% training time cost compared to LLM-based models (i.e., LISA). We also record the evaluation time on RefCOCO/testA (750 images with 5556 referring instances) per image. Our EVF-SAM can reduce 63% to 83% inference time cost, too.
4.5 Ablation Study
In this section, we conduct experiments to investigate the vision-language models for text-prompted SAM and study the effects of the designs of the proposed EVF-SAM. A few ablations are listed in the supplementary. Unless specified, we mainly report the cIoU on testA of RefCOCO.
Table 5: Comparison of effects of different foundation models. The main bottleneck of EVF-SAM lies in the the Multimodal Encoder. Our proposed architecture can perform well on various SAM-like methods.
@@ 4.5 Ablation Study
Multimodal Encoder and fusion methods. In Tab. 4, we explore the effects of different Multimodal Encoders, e.g., CLIP, ViLT [23], and BEIT-3, and fusion methods, e.g., late fusion or early fusion. As shown in Tab. 4, using a text-only encoder in EVF-SAM obtains limited segmentation performance on RefCOCO. Using Multimodal Encoders with both image and text inputs remarkably improves 4.5 cIoU, 4.6 cIoU, 4.0 cIoU, 1.0 cIoU, and 4.5 cIoU for CLIP-Large (OpenAI  [^3] OpenAI: https://github.com/openai/CLIP), CLIP-Large (OpenCLIP  [^4] OpenCLIP: https://github.com/mlfoundations/open_clip), CLIP-Huge (OpenCLIP), ViLT, and BEIT-3, respectively. It demonstrates the superiority of using multimodal prompts (text and input image) and showcases that the image embeddings will also provide useful guidance for SAM to segment objects accurately. We further evaluate the effects of early fusion on ViLT and BEIT-3, which adopts modality fusions in all self-attention layers. Specifically, we adopt two settings for BEIT-3 to analyze, e.g., fusions among former 12 layers ($L_{1}\sim L_{12}$), and fusions among all layers ($L_{1}\sim L_{24}$). Tab. 4 indicates that BEIT-3 with early fusion (fusing former 12 layers or fusing all 24 layers) significantly improves compared to late fusion or using text only. In addition, ViLT with early fusion also achieves 11.1 cIoU improvements compared to the baseline with text-only prompts, showing the effectiveness of early fusion and multimodal inputs for prompting SAM. Therefore, Tab. 4 demonstrates that (1) Multimodal Encoder with the input image and text and (2) early fusions between the image and text encoder are much effective for text-prompted SAM.
Table 6: Ablation on trainable modules. We mainly evaluate the effects of fine-tuning or freezing the Multimodal Encoder, the prompt encoder and mask decoder of SAM. ‘✓’ denotes trainable, while $\ast$’ denotes frozen.
Multimodal Enc. Prompt Enc. Mask Dec. cIoU
$\ast$ $\ast$ 21.2
$\ast$ $\ast$ 82.9
$\ast$ 83.3
83.7
Ablation on trainable modules. In Tab. 6, we evaluated the effects of fine-tuning (✓) or freezing ($\ast$) modules in the proposed EVF-SAM, i.e., the Multimodal Encoder, the prompt encoder, and the mask decoder. The image encoder of SAM is kept frozen during training. As Tab. 6 shows, fine-tuning the Multimodal Encoder is crucial and it adapts the Multimodal Encoder to encode text and image inputs to multimodal representation for referring image segmentation. Notably, EVF-SAM can achieve competitive results with all modules of SAM kept frozen, and it can be seamlessly regarded as a strong extension for the original SAM, which simultaneously supports text prompts, box prompts and point prompts. Tab. 6 shows fine-tuning the prompt encoder and mask decoder of SAM brings significant improvements.
Effects of Different Foundation Models. In Tab. 5, we explore the effects of using different foundation models in EVF-SAM. For the Multimodal Encoder, we adopt CLIP-Large (only text encoder), ViLT, BEIT-3-Large, and BEIT-3-Base. We also modify EVF-SAM with Efficient-SAM [64] and SAM-2 [48] to formulate a lighter or better version. As shown in Tab. 5, EVF-SAM with BEIT-3-Base brings a severe performance drop which indicates a better Multimodal Encoder leads to better prompts for SAM. Remarkably, Tab. 5 shows a negligible difference between Efficient-SAM-S and SAM-H in EVF-SAM, which indicates that EVF-SAM performs well for different SAM variants. In addition, it also provides insights about designing text-prompted SAMs for future research, e.g., developing a larger and better Multimodal Encoder is more important to empower SAM with text-prompted abilities.